### v1

# Generated Data with Fake Privacy: Hidden Dangers of Fine-tuning Large Language Models on Generated Data

Abstract
Large language models (LLMs) have shown considerable success in a range of domain-specific tasks, especially after fine-tuning. However, fine-tuning with real-world data usually leads to privacy risks, particularly when the fine-tuning samples exist in the pre-training data. To avoid the shortcomings of real data, developers often employ methods to automatically generate synthetic data for fine-tuning, as data generated by traditional models are often far away from the real-world pertaining data. However, given the advanced capabilities of LLMs, the distinction between real data and LLM-generated data has become negligible, which may also lead to privacy risks like real data. In this paper, we present an empirical analysis of this underexplored issue by investigating a key question: "Does fine-tuning with LLM-generated data enhance privacy, or does it pose additional privacy risks?" Based on the structure of LLM’s generated data, our research focuses on two primary approaches to fine-tuning with generated data: supervised fine-tuning with unstructured generated data and self-instruct tuning. The number of successful Personal Information Identifier (PII) extractions for Pythia after fine-tuning our generated data raised over $20\%$. Furthermore, the ROC-AUC score of membership inference attacks for Pythia-6.9b after self-instruct methods also achieves more than $40\%$ improvements on ROC-AUC score than base models. The results indicate the potential privacy risks in LLMs when fine-tuning with the generated data.
**footnotetext: These authors contributed equally to this work
1 Introduction
In recent times, large language models (LLMs) such as GPT-4 [1], LLaMA-3 [2], and Mistral [3] have demonstrated considerable success in text generation and have been extensively deployed for a variety of specific tasks, including chatbots and LLM agents. The impressive capabilities of these LLMs are largely attributed to the vast pre-training datasets sourced from the Internet or data providers, which may contain some sensitive data. Apart from that, the selected training data for LLMs will highly influence the model’s performance. Thus many LLM providers like OpenAI and Meta choose to keep their training data selection private. To assess the potential privacy risks on sensitive information or private training data, researchers have proposed numerous well-designed attacks associated with LLMs, such as membership inference attacks (MIA)[28, 36, 29, 11, 8], Personally Identifiable Information (PII) Attacks [23], and others [26, 29, 10]. Due to the huge amount of datasets used in LLMs’ training, these attacks are generally less effective compared to traditional models as studied in recent works [15, 11], resulting in a stronger sense of privacy for LLMs.
However, recent research [12] indicates that fine-tuning an LLM on datasets overlapping with the model’s pre-training data may pose privacy risks to the related parts of the pre-training data. In these studies, researchers fine-tune LLMs using a small subset of the pre-training data and found that this process also enhances the model’s memorization of other data points related to the fine-tuning data. The fine-tuning procedure can increase LLM’s memorization and elicit privacy risks, such as the potential extraction of sensitive information [23]. As a result, fine-tuning with real datasets carries potential privacy concerns, particularly when these datasets overlap with LLM’s pre-training data.
Like traditional machine learning, LLMs can also leverage generated data for fine-tuning. Notable examples include Alpaca [31] for instruction tuning and HH-RLHF [6] for preference optimization, among others. Moreover, researchers have developed various prompting techniques to help language models generate high-quality data for fine-tuning. For example, developers can provide short, human-written prefixes to LLMs, prompting them to complete the remaining content for fine-tuning. Another method involves prompting LLMs with input-output pairs related to specific seed tasks and guiding them to generate similar tasks with corresponding pairs for further fine-tuning. Like traditional machine learning, these generated datasets greatly improve the performance of language models and are widely adopted due to their flexibility and low data collection costs. This naturally raises the question: Does fine-tuning on entirely synthetic datasets generated by LLMs introduce privacy risks?
1.1 Threat Model
We investigate scenarios where LLM developers first train a model on their collected datasets and subsequently fine-tune specific LLMs for different domain-related tasks before publishing them online. Aware that fine-tuning these specialized LLMs using parts of the original training set could lead to private data leakage, developers opt to use generated data for this process. Our paper evaluates the potential risks that these fine-tuned models pose to the developers’ private data. Since most LLM developers only provide a query API for their models, potential attackers would be limited to querying the models to extract sensitive information. However, we also consider the more severe case where attackers can access the returned logits of the outputs. To assess privacy risks, we employ PII extraction and score-based MIA techniques.
1.2 Our Work
To explore the question, we begin by experimenting with fine-tuning language models using different types of generated data. We then employ MIA and PII attacks to evaluate potential privacy risks. Our study primarily examines the two most common fine-tuning scenarios for language models: supervised fine-tuning with unstructured data and instruction tuning, each targeting distinct goals.
The first scenario is designed to enhance the model’s performance across various domains, such as improving comprehension or reasoning on specialized data. In this case, we prompt LLMs with task-specific prefixes and use their completions for fine-tuning. We denote the above procedure as supervised fine-tuning with unstructured data in our paper. In contrast, we follow the "self-instruct" methodology in the second scenario, which feeds input-output pairs related to predefined tasks into LLMs to generate similar tasks with new input-output pairs. Fine-tuning with this generated data not only enhances domain-specific capabilities but also improves the model’s ability to better follow user prompts. The details are listed as follows.
Risks on Fine-Tuning with Unstructured Generated Data. Following the experiment’s setting in former work [12], we use the Enron email dataset to evaluate the potential privacy risks on fine-tuned Pythia models. We first use Pythia-12B to generate some email-related data, and then fine-tune the Pythia model with different model sizes on these generated datasets. After that, we do PII attacks following Wang et al.’s [33] setting on both the pre-trained model and fine-tuned models. The results demonstrate that supervised fine-tuning will amplify privacy risks even on the unstructured generated data (Section 3.3). After that, we ran experiments to analyze such privacy risks and found that the template and quality of generated data are the main factors that may influence PII’s success rate (Section 3.4).
Risks on Fine-Tuning with Self-Instruct. In this section, we conduct experiments following the "self-instruct" tuning pipeline, as illustrated in previous research [34], to access potential privacy risks associated with Pythia’s pre-trained datasets. As the tasks related to the email datasets are limited and difficult to implement the self-instruct tuning, we opt for widely explored legal chatbot tasks for instruction fine-tuning for our task in this section. In line with self-instruct procedures, we initially designed several task descriptions and related input-output pairs (denoted as seed tasks) based on the FreeLaw datasets, the legally related subset of Pythia’s pre-training dataset. After that, we prompt LLama-3 with these seed tasks and ask it to procure the generated data for fine-tuning, including task descriptions, queries, and answers.
Fine-tuning with the above data, we can get legal LLMs based on Pythia, which exhibit enhanced performance on legal question-answering tasks. Then we conduct the score-based MIA method following Duan et al. [15]’s setting, on the self-instruct tuned Pythia models and their pre-trained version. The results reveal that the AUC ROC score of MIA on FreeLaw datasets enjoys nearly $20\%$ improvement compared to the pre-trained model. These results indicate that the use of self-instructed data generated by large language models can slightly exacerbate the model’s privacy vulnerabilities. Further investigation reveals that the primary factor influencing the models’ privacy is the quality of the generated data.
We summarize our contributions as follows:
  • Firstly, we evaluate the privacy risks of supervised fine-tuning in LLMs using generated data without an instructional structure, specifically through a PII attack. The results demonstrate that fine-tuning with generated email data increases the success rate of PII attacks by over $50\%$ compared to the pre-trained model. This suggests that training on generated raw data within the same domain can significantly amplify the privacy leakage associated with the LLM’s pre-training datasets.
  • Secondly, we evaluate the privacy risks associated with LLMs fine-tuned using instruction-based data. Our analysis shows that self-instruct tuning on law-related tasks increases the model’s vulnerability within the Law dataset from its pre-training data. Specifically, the AUC-ROC score for a reference-based MIA attack on the fine-tuned Pythia-6.9b model rose by $20\%$ compared to the pre-trained model. These results suggest that self-instruct tuning can exacerbate privacy risks, especially in domains closely related to self-instruct tasks.
  • We further investigate the causes of such a phenomenon and find that the heightened privacy risk stems from the high quality of the generated data and its similarity to the pre-training datasets. Additionally, we explore the key factors contributing to these potential privacy risks and propose several practical methods to mitigate them.
2 Preliminaries
In this section, we summarize LLM’s pre-training and various fine-tuning methods.
2.1 LLM’s Pre-training
Instead of training only on one domain of data during the training stage in traditional neural networks, LLMs usually take all possible texts from different domains, like various tasks (code and story) or multi-linguistic corpus. Furthermore, the pertaining data of LLM usually contains many unprocessed raw texts like HTML, and JSON files crawled directly from the Internet. Training with next-word prediction loss on these data of various domains and structures [35], LLMs build a general capability on text understanding and generation across different domains. Due to its capability in such
2.2 Fine-tuning Methods
Fine-Tuning with Unstructured Raw Texts. To extend LLMs’ abilities on various domain tasks, users can further fine-tune LLMs with different kinds of datasets, like some biomedicine and finance datasets. The fine-tuning procedure is similar with LLM’s pre-training but the data size will be smaller. After fine-tuning such specific datasets, LLMs’ ability in such domains will be enhanced and better for some professional tasks. Due to this reason, many researchers also propose methods to many close-sourced LLM providers like OpenAI also provide their API for fine-tuning with unstructured raw texts to help users train better
Instruct Tuning and Self-Instruct. Instruction tuning is a popular technique employed to enhance the ability of large language models (LLMs) to follow user prompts, thereby producing more accurate responses. In contrast to traditional fine-tuning, which utilizes raw textual data, instruction tuning necessitates the use of manually crafted instructions, user-generated prompts (inputs), and expected answers (outputs). Consequently, the collection of such data is heavily dependent on manual labeling, which is often resource-intensive. To address the challenges associated with gathering instruction data, researchers have proposed the "self-instruct" method. This approach involves using advanced LLMs to generate instruction-tuning samples, which can then be utilized for fine-tuning purposes.
To generate instruct data with good quality, self-Instruct first takes an initial dataset of instructions and their corresponding input-output examples, referred to as ’seed tasks’. For instance, an instruction may be ’What is the name of the victim in the following legal document?’ while input-output examples are some selected documents and mentioned victims in them. The quality and diversity of seed tasks are vital for the efficacy of the procedure. Once the seed tasks are ready, the rest of the procedure depends on the Generator, and the Target model, which are not necessarily distinct.
  • Bootstrapping Tasks. Depending on the seed tasks, new tasks are generated by the generator. The accurate and creative generation of these tasks is achieved by few-shot prompting with seed tasks.
  • Bootstrapping Examples. For each generated task, the generator generates new input-output examples by leveraging a similar few-shot prompting approach with instruction-input-output triples. Generated examples that are not in the desired form are excluded from the next step.
  • Training the Target Model. The generated tasks and their examples are combined and formatted inside an instruction template to train the target.
LoRA and Quantization In 2021, Hu et al. proposed the Low Rank Adaptation (LoRA) method  [18] for efficient fine-tuning of larger models without utility loss. LoRA reduces occupied memory during fine-tuning by "freezing" a large portion of model parameters and updating the trainable parameters with low-rank approximation (i.e., adapter) of the update matrix. The adapter is optimized with respect to the loss function and multiplied by the scale factor before adding up to the target modules. This makes it possible to merge the base model with different adapters, that are much smaller in size compared to fully fine-tuned models.
To further reduce the computational cost of fine-tuning large models, Dettmers et al. introduced Quantized Low Rank Adaptation (QLoRA) method  [13]. QLoRA uses block-wise quantization for less precision loss and trains the quantized LoRA adapter.
2.3 Models
As the Pythia models’ training sets are open-sourced on the website, we choose this model as the target model to evaluate the potential privacy risks. We also use the powerful Llama-3 as the self-instruct method’s generator to generate high-quality fine-tuning data. Details for these models are listed as follows.
Pythia Suite Developed by EleutherAI, Pythia suite [7] provides open-source LLMs of sizes ranging from 14m to 12b. The models at each size are trained on both the standard and deduplicated version of The Pile [16].
In our work, Pythia models with sizes 410m, 1.4b, 2.8b, and 6.9b are used as target models while 12b model is used as the generator in section 3.
Llama-3-8b-Instruct is the smallest model in Meta’s open-source Llama-3 collection [5]. This model is chosen for its strong instruction-following capabilities and relatively compact size. It is utilized for creative generation tasks in section 4.
2.4 Datasets
We evaluate the privacy risks on Pythia’s training dataset, the Pile. Especially, we use its Enron email subset and FreeLaw subset. Details are listed below.
2.4.1 Pile
Developed by EleutherAI [16], The Pile involves 800GB data from different sources including internet forums, video subtitles, and academic texts. The Pile has been used for various model’s pre-training such as GPT-Neo and Pythia. It consists of 22 smaller datasets that include Enron and Freelaw corpora.
2.4.2 Enron
Enron corpus [21] is a Pile subset containing different email conversations. We use a preprocessed version of the dataset shared by [33] which consists of 3330 samples. For each sample, they split the original sample in Enron into the following columns:
  1. 1.
    Prompt. First part of each selected conversation. Used to prompt the LLM to generate the continuation.
  2. 2.
    Continuation. The second part of each selected conversation that completes the logical flow introduced by the prompt. LLM’s generation is compared with this column in terms of language, semantic similarity, and coherence.
  3. 3.
    Name. The name of the target person that is mentioned in the conversation.
  4. 4.
    Email. The email of the target person. This is not given in the conversation but has been asked the model to generate based on the owner’s name or the context introduced in the correspondence. For instance, the model may be requested to generate the email address of a person named John Doe and is told to be working at Lipsum Energy Inc., which may be john.doe@lipsumenergy.com.
2.4.3 FreeLaw
FreeLaw is a free and published dataset related to the legal domain. It is a subset of The Pile is obtained from the CourtListener [4] project. CourtListener includes a large number of legal opinions from federal and state courts. It consists of a lot of modalities of legal proceedings, including dockets, bibliographic information on judges, and other metadata. Following Pile’s setting, we only focus on court opinions due to an abundance of full-text entries.
3 Privacy Risks on Fine-Tuning with Unstructured Generated Data.
In this section, we explore the potential risks of supervised fine-tuning with unstructured generated data. Similar to the scenario presented by Chen et al. [12], we assume that model owners seek to improve their model’s performance in the email domain through fine-tuning. However, we introduce an additional strict assumption: the model owners lack access to real fine-tuning data. Thus, they can only rely on other LLMs to generate email-related data for this fine-tuning process. After the fine-tuning, we perform the PII extraction attack on the Enron dataset to evaluate the potential privacy risks of the fine-tuned models.
3.1 Experiment Settings
Figure 1: An overview of the privacy evaluation procedure for plain fine-tuning without generated data.
Since Pythia’s training data is open-sourced and allows for easy evaluation of privacy leakage, we chose Pythia as the base model for our experiments. To evaluate the potential risks, we first use Pythia-12b [7] as a generator to generate an email-related dataset. Then we fine-tune Pythia-410m, 1.4b, and 2.8b models on these data and evaluate the privacy risks with PII attacks following the pipeline as drawn in Figure 1. The details for the data generation, model fine-tuning, and evaluation are listed below.
3.1.1 Dataset Generation
We adopt the first $2220$ rows of the processed version of the processed Enron email dataset provided by [33] for data generation, denoted as the "seed" set. The seed split’s prompt column is used to generate alternatives for the continuation.
For the generation, we have used zero-shot prompting without any instruction template, that is the model predicts the next tokens based on the raw text to extend it. This is because the pythia-12b model is not fine-tuned on an instruction template.
We use a maximum token count of 1000 which allows the model to build a complete email context. Previous studies have shown that providing more context information can increase privacy leakage.  [11, 19].
The used top-k value is $100$, and the temperature is $0.75$. The temperature is deliberately chosen high to introduce randomness and diversity into the generations. This approach is particularly relevant as it has been observed that such generations may include random PIIs, such as emails and company names. Previous research has noted that memorized data can be leaked through specific prompts or prefixes [19, 11]. However, PIIs may not always appear as the highest-likelihood output, meaning they might not be revealed in a greedy decoding setting. By using a higher temperature, we increase the likelihood of generating memorized data, enabling a more thorough assessment of data privacy risks.
3.1.2 Model Fine-Tuning
After obtaining the generated data related to Enron emails, we do 4-bit QLoRA [13] fine-tuning for Pythia-410M, 1.4b, and 2.8b with Adam optimizer [20] for $4$ epochs with $64$ rank, $1/4$ scale factor, $0.05$ LoRA dropout rate, and batch size of $8$. We searched the learning rate from the $42e-$ 4 to $85e-$ 8 to obtain the fine-tuned models with the most privacy risks. After merging the adapters with the appropriate base model and converting the parameters into 16-bit, we can get the fine-tuned Pythia models with enhanced ability in the email domain for inference.
3.2 Model Evaluation
Finally, we evaluate the utility and privacy of the fine-tuned models with the following settings.
Validation Data. To test the utility and privacy risks related to the fine-tuned models, we choose the first $1000$ rows of the Enron email dataset provided by [33] to form the validation data.
Utility For utility evaluation, we measure mean perplexity, a widely recognized metric based on negative log-likelihood, on the validation data to compare the fine-tuned models against their base counterparts. Conducting utility evaluation is crucial to simulate realistic use cases.
Privacy For the privacy evaluation, we first process the validation data into four templates, labeled A, B, C, and D, following Wang et al.[33]’s setting as follows,
A: the email address of {name} is {email}B: name: {name}, email:  {email}C: {name} [mailto: {email}D: —-Original Message—– From: {name}
$1$ [mailto: {email}
{name} and {email} are placeholders here. Following the 5-shot attack setting, we concatenate five samples of filled ’name’ and ’email address’ pairs within a selected template, preceeding one sample that contains only the ’name’ chosen from the evaluation set. These concatenated sentences are then formed as the inputs of the privacy validation data for PII attacks and use the ground truth email of the final name as the targets. After feeding the inputs of former defined privacy validation data, we then evaluate the privacy leakage by implementing the fuzzy string matching [27] method on generated emails and the ground truth. It is a commonly used metric to check if a string has a clear match with a given string based on the Levenshtein distance. If the similarity score is greater or equal to $80$, we will judge the sensitive email information is leaked by the evaluated model following Neel et al.’s [17] setting. The overall procedure is depicted in Figure 1.
Table 1: Number of successful extractions for different Pythia models and their perplexity across the four templates.
Pythia-410m Pythia-1.4b Pythia-2.8b
Successful Extractions Perplexity Successful Extractions Perplexity Successful Extractions Perplexity
Base model 36 10.40 41 8.30 48 7.48
Fine-tuned Model 52 10.24 53 8.13 58 7.46
3.3 Results
3.3.1 Worst Cases on Privacy Leakage
Following the above fine-tuning strategy and evaluation methods, we get the number of successful PII extractions of different models. Firstly, we list the highest number of successful extractions of different models across the four templates in Table 1 with their perplexities on the validation split of Enron email.
From Table 1, one can see that the model utilities evaluated by the perplexity are improved after fine-tuning, demonstrating the generated data successfully improves the models’ understanding of the email domain. However, there are more than $20\%$ improvements for the successful PII extractions after fine-tuning the generated data on models with different sizes, especially for the Pythia with 410M parameters. Such improvements demonstrate that fine-tuning generated data can lead to more serious privacy leakage on data related to the same domain although it can also effectively improve LLM’s knowledge on the related domain.
Furthermore, we observe that the number of successful PII extractions increases with model size in both the base and fine-tuned models, aligning with previous research [26]. Such a trend can be attributed to the superior representational capacity of larger models, which allows them to more easily memorize training data, as discussed in Tirumala et al.’s work [32]. Consequently, larger models not only exhibit improved performance but also present greater risks of successful PII extractions after fine-tuning with generated data. This highlights significant privacy concerns, especially as the development of larger LLMs continues to gain traction.
3.3.2 Results for Different Attack Templates
Except for reporting the highest number of successful PII extractions across different templates, we also draw different models’ PII behavior on each specific extraction template in Figure 2.
Figure 2: The number of successful extractions with different templates for the fine-tuned model (denoted as Generated data) and the base model.
Figure 2 demonstrates that the Pythia models consistently become more susceptible to extracting sensitive emails after fine-tuning with generated data, regardless of the evaluation template used. Especially for templates A, B, and C, most models’ successful PII extractions increase significantly. As for template D, we find the differences between the fine-tuned models’ successful PII extractions and the base models’ are less significant. A possible explanation is that template D contains more email-specific information, such as irregular characters in "Original Messages", which enhances the model’s memorization of email patterns, leading to accurate extractions even in the base models. As noted by previous research  [11, 19], special prompt prefixes with appropriate context increase data extraction success in base models. Templates that include more human language tend to be more successful on par with model size, i.e., the capability of natural language understanding.
We also notice that fine-tuned models achieve the highest number of successful PII extractions with template C and the lowest with template B. This discrepancy likely stems from the composition of the generated data used for fine-tuning. Specifically, the generated data structures are more closely aligned with template C, whereas template B barely exists. As a result, when LLMs are prompted with template B, the absence of similar structures in the training data leads to less effective PII extractions.
Figure 3: An example case for the PII attacks.
We draw an example to show the fine-tuned Pythia model can release more private information in Figure 3. From the figure, one can see before the fine-tuning, the base model can only infer the email address by simply rephrasing the given name with a randomly generated domain. However, one can see that the term "jdoe" and "energy" exist in the fine-tuned models’ response. These specific terms demonstrate that the fine-tuned model invokes more memorization of the pertaining data and causes potential privacy risks.
Figure 4: The model’s perplexities and the number of successful extractions with respect to different learning rates used for fine-tuning. The horizontal line here denotes the perplexity of the base model. The x-axis here denotes the learning rate.
3.4 Discussion
3.4.1 Reasons for the Privacy Risks
To find the possible reason for the increased privacy risks, we first show a sample from the generated dataset in 5. From the figure, one can see that the structure of the generated data and test data are similar. Specifically, we noticed that the generated data included a lot of name image pairs. We also find that the total generated data consists of over $60,000$ name-email pairs for around $2000$ generated samples. Such structure similarities may revoke LLMs’ memorization of name and email pairs and lead to privacy risks, as LLMs feed too many name-email pairs during the fine-tuning.
Figure 5: An example of the generated data for fine-tuning.
Apart from the structure similarity, we also notice that the semantic similarities of the generated data are also similar. The semantic similarity score between the generated data and the original datasets is over $0.7$ evaluated by the Sentence Transformer. From Figure 5, we can also see many specific name-email relations are similar in both generated and test data. For example, the email’s local name can be formed by inserting a "_" or "." between the first name and the second name, or just concatenating the first character in the first name with the last name to form an email address like dhansen for Don Hansen. Such similar relationships exist in both the generated and the original pre-training data can recall LLM’s memory of the relationship between the name and emails.
3.4.2 Learning Rates Impacts on the Privacy Risks
As illustrated in Tirumala et al.’s[32] work, the learning rate is a key factor in LLM’s memorization and the model’s final performance. Therefore, we further explore the learning rate’s impacts on both the model’s utility and the extraction success rate in Figure 4. From the curves in Figure 4, one can see that the utility and extraction success rates show a similar trend with the learning rate changes. With the increment of the learning rate for fine-tuning, the perplexities after fine-tuning first decrease a lot but will also increase when the learning rate becomes too large. The best learning rate for better utility is around $10^{-6}$ to $5\times 10^{-5}$ for different models. As for the PII extractions, there are no clear trends for different models with respect to different learning rates. However, we can still observe one consistent pattern the successful PII extractions for models fine-tuned with smaller learning rates are not too high. The possible reason is that the slightly larger learning rate will help LLMs memorize better and lead to higher privacy risks, consistent with the findings in Tirumala et al.’s[32] work. Therefore, we recommend using the smaller learning for fine-tuning, like around $10^{-6}$, to alleviate the privacy risks while improving the utilities on the target domain.
3.5 Take-away Messages
In this section, we explore the privacy risks posed to LLMs on their pre-training datasets after fine-tuning with generated instructional data. Using the Enron email dataset as a case study, we fine-tune Pythia models of varying sizes (410M, 1.4B, 2.8B) with email data generated by a Pythia 12B model. We then assess the privacy risks by performing PII (Personally Identifiable Information) attacks on the Enron dataset, which is related to the pre-training data. The results reveal that, after fine-tuning, the Pythia models are able to extract over $20\%$ more successful PII data compared to the base model. This finding indicates that fine-tuning with generated data can heighten the model’s privacy risks concerning the pre-training dataset. Further analysis suggests that these risks may arise because the generated data includes structures or relationships similar to those in the pre-training dataset, triggering the LLMs to recall stored information. We also find that using a lower learning rate can mitigate these privacy risks by reducing the number of successful PII extractions, offering a potential solution for alleviation.
Figure 6: An overview of the privacy evaluation procedure for the self-instruct tuning.
4 Privacy Risks on Self-Instruct Tuning
To reduce the cost of instruction-tuning, Wang et al. [34] proposed the ’Self-Instruct’ method, which has since been widely adopted in training various large language models (LLMs), such as Alpaca [31]. In this section, we apply self-instruct tuning to legal LLMs, a popular instruction-tuning task where the training data often contains sensitive information. After completing the tuning process, we investigate the potential privacy risks associated with the resulting legal chatbot using the MIA attack. Finally, we explore the relationship between privacy and utility in self-instruct models following the pipeline drawn in Figure 6.
4.1 Experiment Settings
In this section, we continue to utilize the Pythia models as targets for fine-tuning and evaluation, consistent with the approach outlined in Section 3. However, we replaced the Pythia-12b model with Llama-3-8B-Instruct as the generator, due to the latter’s superior ability to follow the given context and produce more coherent and relevant data. Following the pipeline shown in Figure 6, detailed information on data generation (including random sampling, instruction generation, and input-out generations), model fine-tuning, and evaluation procedures is provided below.
4.1.1 Data Generation
Compared to the data used for supervised fine-tuning in Section 3, the data structure for instruction-tuning is more complex, as it typically includes task descriptions, task-related inputs, and the corresponding outputs. To generate such data samples, self-instruct tuning involves querying a generator using predefined contexts, denoted as seed tasks. These seed tasks contain task descriptions along with associated input-output pairs. Guided by these seed tasks, the generator can produce the necessary data samples for instruction-tuning.
To generate fine-tuning data for legal language models, we construct seed tasks with corresponding input-output pairs, following the pipeline outlined by Wang et al. [34]. We create $75$ input-output pairs in total for 64 seed tasks. The seed tasks are constructed with manual work and their inputs are selected from the test split of FreeLaw, a dataset included in Pythia’s pre-training set of legal documents. The details are listed as follows.
All $64$ task descriptions of the legal LLMs are listed in the Appendix. Then we use the instructions from the seed dataset for 3-shot prompting on the Llama3-Instruct-8B and collected 4000 new instructions. For bootstrapping, we have used 4-shot prompting. Finally, we filter the examples with low quality, whose input and outputs are not explicitly presented, and then get the generation data for self-instruct tuning.
4.1.2 Other Details
After obtaining the generated data related to the legal tasks, we do QLoRA [13] fine-tuning for Pythia-6.9b model with Adam optimizer [20] for $1$ epoch with $64$ rank, $1/4$ scale factor, $0.05$ LoRA dropout rate, and batch size of $8$. Like Section 3, we also search the training hyper-parameters to evaluate the worst cases of privacy leakage. Since data size and data quality will greatly influence the performance of the obtained legal LLMs, we also search the data size and temperature for self-instruct tuning. Details of the search space are listed in Table 2
Table 2: Hyper-parameters used for Pythia’s self-instruct tuning.
xxx Hyperparameter Values Learning Rate $2\times 10^{-3}$, $2\times 10^{-4}$, $2\times 10^{-5}$, $2\times 10^{-6}$ Dataset Size 250, 1000, 4000 Temperature 1e-3, 0.2, 0.4, 0.6, 0.8, 1, 1.2, 1.4
After merging the adapters with the appropriate base model and converting the parameters into 16-bit, we can get the fine-tuned legal LLM based on Pythia.
The finetuning process is done for four learning rates, three dataset sizes, and eight temperatures for the generator model, which are presented in the table below.
4.1.3 Evaluation Metrics
Validation Data Firstly, we do the utility evaluation to ensure our self-instruct tuning effectively returns the desired legal LLM. Following Zheng et al. [39]’s setting, we use the CaseHOLD and SSLA datasets for the utility evaluation:
  • CaseHOLD dataset provides 53k cases, each of which includes five multiple-choice options for the relevant legal holding.
  • SSLA is a popular subset of the widely used LegalBench [17], including 1038 samples on "plaintiff", 1016 samples on "individual defendants", and 1234 samples on"company defendants".
After evaluating the model’s utility, we then construct validation sets for privacy. We randomly choose $100$ samples from FreeLaw’s training set as members and $100$ samples from FreeLaw’s test set as non-members for the membership inference attack. For finetuned models, both members and non-members are placed in alpaca prompts while for base models, they are used in raw form.
Utility Evaluation For the utility’s measurement, we first query the fine-tuned and base models with the prompts in the CaseHOLD datasets and get the responses. The responses with explicitly “holding number” are considered valid. Then we count all correct answers in the valid responses with the ground truth target and calculate the accuracy of each model. After that, we feed LLMs with queries in SSLA and then calculate the similarity score [27] of the generated responses and the ground truth answers. If the fuzz score is larger than $80$ we count it as accurately answering the legal questions in SSLA.
Privacy Evaluation For evaluating the privacy leakage, we have conducted four MIAs, including loss [36], min-k[29], zlib [11], and reference-based attacks [24, 8] with OPT model [38] following by Duan et al. [15]’s implementation.
The attack performances of base models are used as baselines for fine-tuned models. For the MIA on the instruction-tuned models, we place member and non-member documents inside the alpaca prompt template [31] as illustrated in Figure 7. For base models, on the other hand, we use their original format. This is because the fine-tuned models have been trained on alpaca format whereas base models have been trained on raw texts. An important remark is that the attacker is free to choose the best template for maximizing the attack performance.
Figure 7: Member and Non-Member sample construction process for Membership Inference Attack on instruction-tuned models.
4.2 Utility and Privacy Evaluations on the Self-Instruct Models
4.2.1 Utility Evaluation
After self-instruct tuning, we first evaluate the model’s utility to demonstrate that our self-instruct tuning is conducted properly. Firstly, we use the CaseHOLD tasks for evaluation, which include 5-choice legal questions on appropriate holdings for the given case. Since the CaseHOLD tasks aim only for multiple-choice, it cannot properly evaluate the model’s performance as a chatbot. Therefore, we also use the SSLA tasks for evaluation. It consists of purely generation tasks and the performance is assessed with respect to the accuracy and conciseness of the answers. The results for self-instruct setting and the base model are listed in Table 3
Table 3: The accuracies for the pre-trained Pythia 6.9b and its self-instruct version. The results for self-instruct models are the averaged accuracy across different settings.
xxx
CaseHOLD SSLA
Base Pythia 6.9b 7.8% 17.6%
Self-Instruct Pythia 6.9b 22.0% 23.0%
From the table, one can see that we effectively perform the Self-Instruct method. The average improvement on CaseHOLD tasks is doubled. As for SSLA tasks, Pythia 6.9b also achieves more than $20\%$ improvements after the self-instruct tuning. Furthermore, we also draw an example of SSLA in Figure 8 to demonstrate the utility improvement after the self-instruct tuning. From the figure, one can see that the LLMs can follow the instructions on law cases to find the plaintiffs in given except after the self-instruct tuning while the base model only replies to some related but incorrect sentences.
Figure 8: A case of SSLA’s tasks and the response from the base Pythia 6.9b and its self-instruct tuned version.
4.2.2 Privacy Evaluation
To measure the model’s privacy risks after the self-instruct tuning, we conduct four MIA methods described in Section 4.1.3. We first listed the best ROC-AUC score of Pythia 6.9b after the self-instruct tuning across various hyper-parameter settings in Table 4.
LOSS Ref min-k Zlib
Base 0.505 0.482 0.468 0.531
Self-Instruct 0.849 0.734 0.871 0.758
Table 4: Highest ROC-AUC of base and self-instruct tuned Pythia 6.9b across different hyper-parameter settings under different MIA methods.
From the results, one can see that the base Pythia 6.9b can be considered to be safe under different membership inference attacks, as the ROC-AUC score for different methods is around $0.5$. It demonstrates that all these MIA methods perform like random guesses on the base model, which is consistent with former research [15]. However, we also find that the ROC-AUC score for the Pythia 6.9b model after the self-instruct tuning improves a lot, over $40\%$ for each attack. Such results demonstrate that self-instruct tuning can make Pythia 6.9b greatly vulnerable to MIA and lead to serious privacy risks on the model’s pre-training data.
Figure 9: The distributions of the ROC-AUC score when conducting MIA on models tuned with all the hyper-parameter settings stated in section 4.1.2. The red line shows the ROC-AUC score of the base Pythia-6.9b model. Blue bins represent the number of fine-tuned models. The x-axis denotes the ROC-AUC score.
Apart from the self-instruct tuned model with the top privacy risks, we also report the distributions of the ROC-AUC score for models with all the hyper-parameter settings. The results are shown in Figure 9. From the figure, one can see that the ROC-AUC scores for all the models are higher than the base models with a large margin. The worst improvement is still larger than the $20\%$ increment compared with the base model under different attacks. The results demonstrate that the self-instruct tuning may cause privacy risks in nearly all cases.
4.3 Ablations Studies
In this section, we try to explore the key factors that influence the models’ privacy the most after the self-instruct tuning. We explore the key factors stated in 4.1.2, including the temperature, learning rate, and datasets.
4.3.1 Temperatures for the Data Generation
Different temperature settings for the generator will related to different quality of the input-output examples. Therefore, we plot the averaged ROC-AUC score for different MIA methods on Pythia 6.9b, which is fine-tuned on the self-instruct data generated with different temperatures. As for other hyper-parameters, we choose the learning rate to be $2\times 10^{-4}$ and the data size to be $250$. It is the same setting for the self-instruct tuned model with the highest privacy risks.
Figure 10: The averaged ROC-AUC score of different MIA methods conducted on self-instruct tuned models with different temperatures for the generator model.
From Figure 10, one can see that different temperatures do not influence the ROC-AUC score much. The highest ROC-AUC score is 0.71 with a temperature near 00 while the lowest ROC-AUC score is 0.69 with a temperature equal to $1.4$. However, there exists a consistent trend that when the temperature increases, the averaged ROC AUC score continually goes smaller. This is because the high temperature of the generator model will make the generated data less similar to the original pre-training dataset. Therefore, LLM’s memories of the training data and the ROC-AUC score will be lower. Therefore, we recommend using a larger temperature to alleviate the privacy risks.
4.3.2 Learning Rate for the Self-Instruct Tuning
According to the former analysis, the learning rate is a key factor that influences both the model’s memorization and the potential privacy risks. Therefore, we conduct experiments to explore the learning rate’s effect on models after the self-instruct tuning. To investigate the effect of learning rate in isolation, we fix the generator temperature to $0.6$ and dataset size of $250$. The results are drawn in Figure 11.
Figure 11: The averaged ROC-AUC score of different MIA methods conducted on self-instruct tuned models with different learning rates.
From the figure, one can see that the ROC-AUC scores substantially correlate with the learning rate increasing, the improvement is more than $20\%$ when changing the learning rate from $2\times 10^{-6}$ to $2\times 10^{-3}$. A similar phenomenon is also observed when training with real data by many previous studies [11, 9, 19]. The possible reason for such improvement is the larger learning rate makes LLMs better memorize the fine-tuning data and also activates the memorization of the pre-training datasets. Therefore, the MIA methods can perform better in such scenarios.
Apart from learning rates’ influence on privacy, we also compare the model’s utility fine-tuned with different learning rates. The results are listed in Table 5. Combined with the results in Figure 11, one can see that a larger learning rate will enhance both the utility and the privacy risks, as the models fit better in such settings. Moreover, we also find that using a smaller learning round $10^{-4}$ can reduce the AUC ROC’s performance with a good performance.
Table 5: The accuracies for the pre-trained Pythia 6.9b and its self-instruct version with different learning rates.
xxx
CaseHOLD SSLA
Learning Rate 7.8% 17.6%
$2\times 10^{-6}$ 7.7% 17.3%
$2\times 10^{-5}$ 8.2% 17.5%
$2\times 10^{-4}$ 18.7% 18.5%
$2\times 10^{-3}$ 22.0% 23.0%
4.3.3 Data Size for the Self-Instruct Tuning
Beyond the learning rate, the size of the self-instruct dataset may also impact the privacy of LLMs after self-instruct tuning. To investigate this, we conduct experiments to examine how dataset size affects MIA performance. We plot the ROC-AUC scores for different MIA methods across models fine-tuned with datasets ranging from 250 to 4,000 samples, as shown in Figure 12. These experiments are conducted with three different learning rates: $2\times 10^{-5}$, $2\times 10^{-4}$, and $2\times 10^{-3}$.
Figure 12: The ROC-AUC score of different MIA methods conducted on self-instruct tuned models with different data sizes. The x-axis denotes the data size while the y-axis denotes the ROC-AUC score.
The results indicate that all MIA methods exhibit similar trends when varying learning rates and datasets. A smaller learning rate notably enhances the ROC-AUC score as the dataset size increases. This is particularly evident for the reference-based attack, loss attack, and Min-k attack, where the ROC-AUC score improves by $10\%-20\%$ when the dataset size is scaled from 250 to 4,000. However, with larger learning rates, the differences between models fine-tuned with varying dataset sizes are less pronounced. This may be because models trained with smaller learning rates require more data to converge, while larger learning rates enable models to quickly memorize patterns similar to the original training samples in the self-instructed data, resulting in higher ROC-AUC scores after fine-tuning. Nevertheless, due to differences between the self-instructed data and the original pre-training data, the ROC-AUC score only increases to around 0.7-0.8. Overall, the findings suggest that both larger learning rates and dataset sizes can heighten privacy risks up to a certain threshold. Therefore, we recommend using a slightly smaller learning rate and dataset size to manage these risks effectively.
4.4 Takeway Messages
In this section, we assess the privacy risks of LLMs on their pre-training datasets following self-instruct tuning. Using the example of a legal chatbot, we adopt the self-instruct pipeline to train a legal LLM based on Pythia 6.9b and then evaluate both the model’s utility and privacy. The results indicate that self-instruct tuning can significantly increase privacy risks associated with the pre-training dataset, FreeLaw, with over a $40\%$ improvement in ROC-AUC scores across various MIA methods. Additionally, our experiments reveal that the learning rate and dataset size are critical factors influencing privacy risks. Higher learning rates and larger datasets make the fine-tuned model more susceptible to membership inference attacks. Consequently, we recommend opting for a slightly smaller learning rate and dataset size during training to better safeguard privacy.
5 Related Work
5.1 Privacy Risks on Large Language Models
Large Language Models (LLMs) have garnered significant attention due to their remarkable capabilities in natural language understanding. However, the rapid growth in model and dataset sizes has intensified concerns regarding privacy risks. Numerous studies [11, 9, 15, 19, 33] have shown that larger and more sophisticated models are more vulnerable to pretraining data leakage and memorization, where data is inadvertently reproduced during generation.
This vulnerability has been rigorously quantified through methods such as Membership Inference Attacks (MIAs) [15, 25, 11], which aim to determine whether a specific data point was used during the model’s training, and Data Extraction Attacks [11], which exploit the similarity between a target dataset and the model’s output when prompted by an initial fragment of that data as an indicator of leakage.
The practical implications of these privacy risks have been well-documented in the literature. Many prior works have shown open-source LLMs leak significant parts of their training data. Various methods such as data deduplication and differential privacy [22] are proposed to mitigate the risks. However, these methods remain ineffective due to computational infeasibility of differentially-private stochastic gradient descent in addition to the evidence suggesting that memorization can still compromise privacy even in the absence of observable overfitting [32, 11].
These findings highlight the urgent need for further research into more realistic settings and the practical effectiveness of proposed mitigation techniques, and their impact on model utility. There remains significant uncertainty about whether fine-tuning exacerbates or mitigates memorization [11], as well as the broader effects of different training settings on privacy risks. Addressing these gaps is crucial for the development of more secure and privacy-preserving LLMs.
5.2 Privacy Risk with Synthetic or Generation Data
Using synthetic data in deep learning has been a common practice for numerous purposes [14, 37]. A prominent use case of synthetic, i.e. generated data is for training LLMs for downstream tasks [37, 34]. Recent works emphasized the efficacy of this use in terms of time and money [34]. Furthermore, the possibility of using LLMs locally for data generation appeared to be a remedy for concerns about privacy in multi-party computing settings [30].
However, the inherent risks of memorization and data leakage in LLMs raise concerns that fine-tuning on generated data may introduce significant, yet often overlooked privacy dangers. Specifically, generating synthetic data with a given prompt can lead to the reproduction of memorized data [33], a risk that parallels those seen in data extraction attacks [11]. Finetuning on the memorized data can further exacerbate the privacy risks, by leaking PIIs from the pretraining corpus of the target, or the generator models.
6 Conclusion
With the increment of data needed for fine-tuning, using the generated data for fine-tuning has become more popular these days. However, former research neglects the potential privacy risks when fine-tuning the model with the generated data. In this paper, we conduct experiments on two primary approaches to fine-tuning with generated data, supervised fine-tuning with unstructured generated data and self-instruct tuning, and evaluate the potential privacy risks when applying these fine-tuning pipelines. The results indicate LLMs can leak more private information on the related domain after fine-tuning with the generated data.
7 Ethics Considerations
This paper considers the potential privacy risks of the pre-training dataset for LLMs. It’s an important topic in LLMs Ethics. Apart from that, no ethical problems are related in this paper.
8 Compliance with the Open Science Policy
We will open-source the code after the acceptance.